OpenStudy (anonymous):

A set of numbers is said to be closed under an operation if the result of combining any two numbers in the set results in a number that is also in the set. Decide whether or not each set is closed under the operation. a. {positive integers]; division b. {odd integets}; x c. {odd integers}; + d. {integers ending in 4 or 6}; x

OpenStudy (swissgirl):

Positive integers divided by positive integers will only give you positive integers

OpenStudy (swissgirl):

positive divided by positive will result in positive positive divided by negative will result in negative negative divided by negative willl result in a negative

zepdrix (zepdrix):

For a. lemme give you an example: \(\Large\rm \dfrac{5}{3}\) There is a positive integer, dividing a positive integer. Is the result a positive `integer`?

OpenStudy (anonymous):

Yeah.

zepdrix (zepdrix):

No, 5/3 is not an integer.

OpenStudy (anonymous):

That's a ratio, right?

zepdrix (zepdrix):

Lemme pick some better numbers ^^ hehe How bout 10 divided by 4? `positive integer` divided by `positive integer`, but the result is `2.5`, which is NOT an integer, right? :o

OpenStudy (anonymous):

Yeah...

zepdrix (zepdrix):

We have `left` the set of positive integers by using division. So we're not closed under division for the first set, yah? :)

OpenStudy (anonymous):

Oh yah. So for b, the set is closed, right?

zepdrix (zepdrix):

For b. and c. you need to remember a little something about even and odd numbers. `odd` times `odd` = ? `odd` plus `odd` = ? 3 x 7 = 21 5 x 5 = 25 9 x 3 = 27 `odd` times `odd` = `odd` Yes, good :)

OpenStudy (anonymous):

odd plus odd = positive 1+1=2 3+3=6 so is C, not closed?

zepdrix (zepdrix):

Correct. Because {odd integer} + {odd integer} =/= {odd integer} Open.

OpenStudy (anonymous):

Is D closed too?

zepdrix (zepdrix):

lol are you just guessing? :) How do you conclude that for D?

OpenStudy (anonymous):

Haha, kind of :3 4x4=16 6x6=36 4x6=24

zepdrix (zepdrix):

Ah good good. You really only need to check 3 cases, yah? I'm gonna use a sloppy notation a sec. So if you have a number ending in 4, ...4 and multiply it by a number ending in 4, ...4 x ...4 = .....16 It ends in a 6, regardless of what the other numbers are. That works out for the other cases as well. Yay good job \c:/

OpenStudy (anonymous):

^_^ Thank youu [: